I don't ned answers, I need to know how to come up with the answers

When the price goes down, the quantity demanded goes up. The price elasticity of demand measures ______

2. The price of gasoline rises 5% and the quantity of gasoline purchased falls 1%. The price elasticity of demand is equal to _____, and demand is described as _____.

3. The ratio of the percentage change in quantity demanded to the percentage change in price is the _____ elasticity of demand.

4. The price elasticity of demand can be found by _______.

5. If the estimated price elasticity of demand for foreign travel is 4: A) a 20% decrease in the price of foreign travel will increase quantity demanded by 80%. B) the demand for foreign travel is inelastic. C) a 10% increase in the price of foreign travel will increase quantity demanded by 40%. D) a 20% increase in the price of foreign travel will increase quantity demanded by 80%.
Answer #1

1. When the price goes down ,the quantity demanded goes up . The price elasticity of demand measures the responsiveness of the quantity change in relation to the price change.

2. The price of gasoline rises 5% and the quantity of gasoline purchases falls 1% . The price elasticity of demand is equal to 1/5 = 0.2 , and demand is described inelastic because it is less than 1.

3. The ratio of the percentage change in quantity demanded to the percentage change in price is price elasticity of demand.

4. The price of elasticity of demand can be found by comparing the percentage change in quantity demanded to the percentage change in price.

5. If the estimated price elasticity of demand for foreign travel is 4 , this implies that a 20% increase in the price of foreign travel will increase the quantity demanded by 80% because then , Ed= 80/20= 4.Hence,option(A) is correct.